# TopER: Topological Embeddings in Graph Representation Learning

### v2

@@ 1 Introduction
Graphs are a fundamental data structure utilized extensively to model complex interactions within various domains, such as social networks (Leskovec et al.,, 2008), molecular structures (You et al.,, 2018), and transportation systems (Duan et al.,, 2022). Their inherent flexibility, however, introduces significant challenges when applied to machine learning tasks, primarily due to their irregular and high-dimensional nature. Due to the fact that graph data lacks inherent ordering and consistent dimensionality, traditional machine learning methods—designed for data in vector spaces—struggle with it.
@@
Graph Neural Networks (GNNs) have emerged as the state-of-the-art models for tackling graph machine learning tasks due to their ability to learn effectively from graph-structured data. In the predominant paradigm of message-passing GNNs, the process begins by generating node embeddings. These embeddings can then be used in tasks such as node classification or link prediction. However, for graph-related tasks, such as molecular property prediction, the embeddings must be aggregated through a pooling layer to form graph-level representations. This method is computationally intensive, largely because the generation and management of node embeddings as intermediate steps substantially increase the overall computational burden. Ideally, an approach would allow for the direct creation of graph embeddings, circumventing the need to generate node-level representations first. Furthermore, these graph embeddings must be both low-dimensional and interpretable to maximize their practical utility and efficiency in various applications.
@@
Topological Data Analysis (TDA) is well-suited for directly constructing graph representations without costly node embeddings. Topology studies the shape of data, and TDA primarily focuses on the qualitative properties of space, such as continuity and connectivity (Coskunuzer and Akçora,, 2024). A particularly effective technique in TDA is Persistent Homology (PH), which tracks topological features—like connected components and cycles—across various scales via a process known as filtration. Filtration is adept at revealing both local and global structures within graphs. It proves exceptionally useful for comparing graphs of different sizes that maintain the same inherent structure, which may suggest similar properties in graph datasets. For instance, similar substructures in protein interaction networks across different species may indicate comparable biological functions. By focusing on data shape, PH proves invaluable in graph tasks that benefit from a graph-centric approach, offering insights that might not be as apparent when focusing on individual node analysis. This shift from a node-centric to a graph-centric perspective can dramatically improve the understanding and application of graph data in fields like bioinformatics and network analysis. However, the utility of Persistent Homology is limited by the high computational demands involved in extracting topological features during the filtration process, mainly due to its cubic time complexity (Otter et al.,, 2017). This constraint reduces its practicality for large-scale graphs and has restricted the broader integration of PH in graph representation learning.
@@
With this work, we take a significant step forward in addressing the challenges of topological graph representation learning and introduce Topological Evolution Rate (TopER). This novel approach refines the Persistent Homology process to efficiently capture graph substructures, thereby mitigating the significant computational demands of calculating complex topological features. As graph representation learning aligns naturally with Topological Data Analysis, TopER excels in graph clustering and classification tasks where it achieves the best rank in experiments. Furthermore, simplifying graph data into a low dimensional space, TopER creates intuitive visualizations that reveal clusters, outliers, and other essential topological features, as demonstrated in Figure 1. As a result, TopER merges interpretability with efficiency in graph representation learning, providing an ideal balance that can scale to large graphs.
To our knowledge, TopER is the first topology-based graph representation learning method that can create low-dimensional, efficient, and scalable graph representations.
Figure 1: TopER Visualizations. In the figures above, each point represents an individual graph. On the left, TopER is applied to three benchmark compound datasets using closeness sublevel filtration. The middle panel zooms in on the red point cloud from the left, demonstrating TopER’s effectiveness in distinguishing between classes within the MUTAG dataset. On the right, a TopER visualization for the IMDB-B dataset is displayed.
@@
  • New Graph Embedding Method: We introduce TopER, a compact and computationally feasible graph representation designed to capture the evolution of graph substructures.
  • Interpretable Visualizations: TopER generates interpretable, low-dimensional embeddings, enabling clear visualization of clusters and outliers. It excels in providing insights within individual graph datasets and across multiple datasets, facilitating comparative analysis.
  • Enhanced Use of Persistent Homology: TopER optimizes the filtration process in PH, offering computational efficiency and compact outputs, while boosting the performance of traditional PH methods.
  • Theoretical Stability: We establish theoretical stability guarantees for TopER, ensuring that the embeddings are reliable and robust across various filtration functions.
@@
  • Competitive Performance: Our extensive experiments on benchmark datasets demonstrate that TopER delivers consistently competitive or superior performance in clustering and classification tasks compared to state-of-the-art models.
2 Related Work
Graph Representation Learning. Graph representation learning is a dynamic subfield of machine learning, focusing on transforming graph data into efficient, low-dimensional vector representations that encapsulate essential features of the data (Hamilton,, 2020; Gao et al.,, 2019). These representations facilitate a deeper analytical understanding of graphs, which is critical for various applications such as molecular graph property prediction (Dong et al.,, 2019).
@@ 2 Related Work
Graph Neural Networks. Graph Neural Networks have revolutionized the analysis of graph data, drawing parallels with the success of Convolutional Neural Networks in image processing (Errica et al.,, 2020). GNNs utilize spectral and spatial approaches to graph convolutions based on the graph Laplacian and direct graph convolutions, respectively (Bruna et al.,, 2014; Defferrard et al.,, 2016; Kipf and Welling,, 2017). Despite their success, GNNs often suffer from issues like over-smoothing and lack transparency, making them less ideal for applications requiring interpretability (Günnemann,, 2022).
@@
TDA in Graph Representation Learning. Topological Data Analysis provides a robust and computationally efficient framework to address the interpretability and over-smoothing issues present in GNNs (Aktas et al.,, 2019). Persistent Homology, a key technique in TDA, has been applied successfully to graph data, demonstrating potential to match or even exceed the performance of traditional methods in classification and clustering tasks (Hensel et al.,, 2021; Demir et al.,, 2022; Hiraoka et al.,, 2024; Immonen et al.,, 2024; Chen et al., 2024a, ; Loiseaux et al.,, 2024). However, the computational intensity of PH limits its scalability (Hofer et al.,, 2019; Zhao et al.,, 2020; Akcora et al.,, 2022).
@@
Graph Embeddings and Visualization. Graph embedding techniques, including spectral methods, random walk-based approaches, and deep learning-based models, transform graph data into vector representations to support tasks like visualization and machine learning (Cai et al.,, 2018; Goyal and Ferrara,, 2018; Xu,, 2021). Approaches such as Laplacian Eigenmaps and DeepWalk have been particularly effective in revealing clusters within graphs (Belkin and Niyogi,, 2001; Perozzi et al.,, 2014). However, these methods are predominantly applied to visualize a single graph in node classification tasks, focusing on cluster identification (Wang et al.,, 2016; Mavromatis and Karypis,, 2020; Tsitsulin et al.,, 2023). Furthermore, they often overlook domain-specific information, which can limit their effectiveness in more specialized applications (Jin and Zafarani,, 2020).
@@
3 Background
TDA has emerged as a promising approach in graph representation learning (Aktas et al.,, 2019). Among its various techniques, persistence homology (PH) is notable for its robust capacity to quantify and monitor topological features across different scales.
@@ 3.1 Persistent Homology for Graphs
Persistent Homology applies algebraic topology to reveal hidden shape patterns within data, capturing these insights by tracking the evolution of topological features like components, loops, and cavities at varying resolutions (Coskunuzer and Akçora,, 2024). The PH process involves constructing filtrations, obtaining persistence diagrams, and vectorizations. Our model, however, primarily utilizes the filtration step, reformulating the evolution information in a novel manner.
In the crucial filtration step, PH decomposes a graph $\mathcal{G}$ into a nested sequence of subgraphs $\mathcal{G}_{1}\subseteq\mathcal{G}_{2}\subseteq\ldots\subseteq\mathcal{G}_{n}=\mathcal{G}$. For each $\mathcal{G}_{i}$, an abstract simplicial complex $\widehat{\mathcal{G}}_{i}$ is defined, forming a filtration of simplicial complexes. Clique complexes are typical choices, where each $(k+1)$-complete subgraph in $\mathcal{G}$ corresponds to a $k$-simplex (Aktas et al.,, 2019). To obtain effective filtrations, utilizing relevant filtration functions is essential.
Filtration Functions. Filtration functions are essential in PH. For a given graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, a common approach is to define a node filtration function $f:\mathcal{V}\to\mathbb{R}$, which establishes a hierarchy among the nodes. By selecting a monotone increasing set of thresholds =$\mathcal{I}=\{\epsilon_{i}\}_{i=1}^{n}$, this method generates subgraphs $\mathcal{G}_{i}=(\mathcal{V}_{i},\mathcal{E}_{i})$ where $\mathcal{V}_{i}=\{v\in\mathcal{V}\mid f(v)\leq\epsilon_{i}\}$ and $\mathcal{E}_{i}$ is the set of edges in $\mathcal{E}$ with endpoints in $\mathcal{V}_{i}$. This is called a sublevel filtration induced by $f$ (See Figure 2). Similarly, superlevel filtrations can be constructed by defining $\mathcal{V}_{i}=\{v\in\mathcal{V}\mid f(v)\geq\epsilon_{i}\}$ for decreasing thresholds (Aktas et al.,, 2019).
Similarly, one can use edge filtration functions $g:\mathbb{E}\to\mathbb{R}$ to define such a filtration. Similarly, by defining $\mathcal{E}_{i}=\{e_{jk}\in\mathcal{E}\mid g(e_{jk})\leq\epsilon_{i}\}$, and $\mathcal{V}_{i}$ as the all endpoints of $\mathcal{E}_{i}$, one can define a nested sequence $\{\mathcal{G}_{i}\}_{i=1}^{n}$. Especially, for weighted graphs, this method is highly preferable as weights naturally defines an edge filtration function. The common node filtration functions are degree, betweenness, centrality, heat kernel signatures, and node functions coming from the domain of the datasets (e.g., atomic number for molecular graphs). Common edge filtration functions are Ollivier and Forman Ricci curvatures, and edge weights (e.g. transaction amounts for financial networks).
@@
We developed a new filtration function, Popularity, in addition to existing approaches, to enhance the topological representation of graph-structured data in our studies. Popularity is similar to the degree function (Newman,, 2003), but also considers the average degree of neighboring nodes of a given node $v$. This function enriches node filtration by incorporating broader neighborhood information, potentially revealing deeper insights into graph structure. Mathematically, for each node $v$ in the graph, we define the popularity function as: $\mathcal{P}(v)=\deg(v)+\frac{\sum_{u\in\mathcal{N}(v)}\deg(u)}{|\mathcal{N}(v)|}$, where $\deg(v)$ is the degree of node $v$ and $\mathcal{N}(v)$ denotes the set of nodes adjacent to $v$. This function incorporates $2$-neighborhood information by giving more weight to high-degree (popular) neighbors. Popularity can be considered an improved version of the degree function, as it also accounts for the popularity of a node’s neighbors. The intuition is that if degree represents the number of friends, popularity accounts for the popular friends (node neighbors with high degree) with greater weight.
@@ 4 TopER: Topological Evolution Rate
TopER innovatively extends the concept of filtration used in Persistent Homology to track graph substructures more efficiently. This reformulation reduces the computational overhead typically required for detailed topological feature extraction. Unlike traditional Persistent Homology, which extracts costly topological features, TopER summarizes the filtration process through two key parameters derived via regression: filtration sequences and evolution.
Filtration sequences. We first decompose a graph $\mathcal{G}$ into a nested sequence of subgraphs (filtration graphs) $\mathcal{G}_{1}\subseteq\mathcal{G}_{2}\ldots\subseteq\mathcal{G}_{n}=\mathcal{G}$ by using a filtration function, such as node degree or closeness. Let $\mathcal{G}_{i}\subset\mathcal{G}$, $\mathcal{V}_{i}$ represent nodes in $\mathcal{G}_{i}$ and $\mathcal{E}_{i}$ represent the edges. Next, we compute $x_{i}=|\mathcal{V}_{i}|$ as the count of nodes, and $y_{i}=|\mathcal{E}_{i}|$ as the count of edges. Then, for each filtration graph $\mathcal{G}_{i}$, we obtain the pair $(x_{i},y_{i})\in\mathbb{R}^{2}$, which creates two monotone sequences, $x_{1}\leq x_{2}\leq\dots\leq x_{n}$ and $y_{1}\leq y_{2}\leq\dots\leq y_{n}$. Hence, TopER yields two ordered sets $\mathcal{X},\mathcal{Y}$ describing the evolution of the filtration graphs $\mathcal{G}_{1}\subseteq\ldots\subseteq\mathcal{G}_{n}=\mathcal{G}$, $\mathcal{X}=(x_{1},x_{2},\dots,x_{n})$ and $\mathcal{Y}=(y_{1},y_{2},\dots,y_{n})$. Here, $n$ corresponds to number of thresholds $\{\epsilon_{i}\}_{i=1}^{n}$ used in the filtration step.
@@
Consider the top row of Figure 2 where we have three filtration graphs (i.e., $n$=3); we have $\mathcal{X}=(2,3,7)$ for node counts and $\mathcal{Y}=(0,2,6)$ for edge counts.
Evolution. In the next step, PH would typically compute topological features on each filtration and create a persistence diagram to summarize the features. Not only it is costly, but the approach would require efforts to vectorize the persistence diagrams. We circumvent this computationally costly step and analyze how the number of edges $\{y_{i}\}$ relates to the number of nodes $\{x_{i}\}$ throughout the filtration sequence. We use line fitting to characterize this relationship as follows.
Simple linear regression, often applied through the least squares method (James et al.,, 2013), is a standard approach in regression analysis for fitting a linear equation to a set of data points $\{(x_{i},y_{i})\}\subset\mathbb{R}^{2}$. This method calculates the line $\mathcal{L}(x)=a+bx$ that best fits the data by minimizing the loss function $\mathrm{E}=\sum^{N}_{i=1}[\mathcal{L}(x_{i})-y_{i}]^{2}$. The coefficients $(a,b)$, obtained from this regression on the filtration sequences, effectively describe the graph’s structure during filtration (see the descriptor step in Figure 3). Algorithm 1 unifies the filtration and evolution steps.
Algorithm 1 TopER: Topological Evolution Rate
1:Input: Graph $\mathcal{G}$, Filtration function $f:\mathcal{V}\to\mathbb{R}$, Threshold set =$\mathcal{I}=\{\epsilon_{i}\}_{i=0}^{n}$
2:Output: TopER vector $\mathcal{T}(\mathcal{G},f,\mathcal{I})$
3: Initialize lists $\mathcal{X}=[]$, $\mathcal{Y}=[]$
4:for $i=1$ to $n$ do
5:     $\mathcal{G}_{i}\leftarrow$ Induced subgraph of $\mathcal{G}$ where $\mathcal{V}_{i}\subseteq f^{-1}([\epsilon_{0},\epsilon_{i}])$
6:     $x_{i}\leftarrow|\mathcal{V}_{i}|$
7:     $y_{i}\leftarrow|\mathcal{E}_{i}|$
8:     Append $x_{i}$ to $\mathcal{X}$
9:     Append $y_{i}$ to $\mathcal{Y}$
10:end for
11:Fit a line $\mathcal{L}(x)=a+bx$ to pairs $(x_{i},y_{i})$ from lists $\mathcal{X}$ and $\mathcal{Y}$ using least squares
12:Extract coefficients $a$ and $b$
13:Return $(a,b)$ as the TopER vector $\mathcal{T}(\mathcal{G},f,\mathcal{I})$
@@
Definition 1 (Topological Evolution Rate (TopER)).
Let $f:\mathcal{V}\to\mathbb{R}$ be a filtration function on graph $\mathcal{G}$ and =$\mathcal{I}=\{\epsilon_{i}\}_{i=1}^{n}$ be the threshold set. Let $\mathcal{G}_{i}=(\mathcal{V}_{i},\mathcal{E}_{i})$ be the induced filtration. Let $x_{i}=|\mathcal{V}_{i}|$ and $y_{i}=|\mathcal{E}_{i}|$. Let $\mathcal{L}(x)=a+bx$ be the best fitting line to $\{(x_{i},y_{i})\}_{i=1}^{n}$. Then, we define theTopER vector of $\mathcal{G}$ with respect to $f$ as $\rm{TE}$$(\mathcal{G},f,\mathcal{I})=(a,b)$. We call $a$ the pivot and $b$ the growth of $\mathcal{G}$.
In practice, line fitting may not always be a linear choice and requires multiple controls, which we cannot report here due to space limitations. We discuss these choices and linear/polynomial fit options in Appendix B. We also show visual samples of graph evolution rates in Appendix Figure 6. Furthermore, we define and visualize the most fundamental evolution patterns in Appendix Figure 7.

### v3

@@ 1 Introduction
Graphs are a fundamental data structure utilized extensively to model complex interactions within various domains, such as social networks (LBKT, 08), molecular structures (YLY+, 18), and transportation systems (DCS+, 22). Their inherent flexibility, however, introduces significant challenges when applied to machine learning (ML) tasks, primarily due to their irregular and high-dimensional nature. Graph data lacks inherent ordering and consistent dimensionality, making it challenging for traditional ML methods designed for vector spaces.
@@
Graph Neural Networks (GNNs) have emerged as the state-of-the-art models for tackling graph machine learning tasks due to their ability to learn effectively from graph-structured data (KW, 17). In the predominant paradigm of message-passing GNNs, the process begins by generating node embeddings (BHG+, 21). These embeddings can then be used in tasks such as node classification or link prediction. However, for graph-related tasks, such as molecular property prediction, the embeddings must be aggregated through a pooling layer to form graph-level representations (WKK+, 20). This method is computationally intensive, mainly because generating and managing node embeddings as intermediate steps substantially increase the overall computational burden. Ideally, an approach would allow for the direct creation of graph embeddings, circumventing the need to generate node-level representations first. Furthermore, these graph embeddings must be both low-dimensional and interpretable to maximize their practical utility and efficiency in various applications.
@@
Topological Data Analysis (TDA) is well-suited for directly constructing graph representations without costly node embeddings. Topology studies the shape of data, and TDA primarily focuses on the qualitative properties of space, such as continuity and connectivity (CA, 24). A particularly effective technique in TDA is Persistent Homology (PH), which tracks topological features, like connected components and cycles, across various scales via a process known as filtration. Filtration is adept at revealing both local and global structures within graphs. It proves exceptionally useful for comparing graphs of different sizes that maintain the same inherent structure, which may suggest similar properties in graph datasets. For instance, similar substructures in protein interaction networks across different species may indicate comparable biological functions. By focusing on data shape, PH proves invaluable in graph tasks that benefit from a graph-centric approach, offering insights that might not be as apparent when focusing on individual node analysis. This shift from a node-centric to a graph-centric perspective can dramatically improve the understanding and application of graph data in fields like bioinformatics and network analysis. However, the utility of Persistent Homology is limited by the high computational demands involved in extracting topological features during the filtration process, mainly due to its cubic time complexity (OPT+, 17). This constraint reduces its practicality for large-scale graphs and has restricted the broader integration of PH in graph representation learning.
@@
With this work, we take a significant step forward in addressing the challenges of topological graph representation learning and introduce Topological Evolution Rate (TopER). This novel approach refines the Persistent Homology process to efficiently capture graph substructures, thereby mitigating the significant computational demands of calculating complex topological features. As graph representation learning aligns naturally with Topological Data Analysis, TopER excels in graph clustering and classification tasks, where it achieves the best rank in experiments. Furthermore, simplifying graph data into a low-dimensional space, TopER creates intuitive visualizations that reveal clusters, outliers, and other essential topological features, as demonstrated in Figure˜1. As a result, TopER merges interpretability with efficiency in graph representation learning, providing an ideal balance that can scale to large graphs. To our knowledge, TopER is the first topology-based graph representation method that can create low-dimensional, efficient, and scalable graph representations.
@@
(a) 3 DATASETS
(b) MUTAG
(c) IMDB-B
Figure 1: TopER Visualizations. Each data point represents an individual graph. On the left, TopER is applied to three benchmark compound datasets using closeness sublevel filtration. The middle panel zooms in on the red point cloud from the left, demonstrating TopER’s effectiveness in distinguishing between classes within the MUTAG dataset. On the right, a TopER visualization for the IMDB-B dataset is displayed.
@@
  • New Graph Representation: We introduce TopER, a principled and computationally feasible graph representation designed to capture the structural evolution of a graph.
  • Topology-Inspired Design: Rather than computing persistent diagrams or Betti numbers, TopER directly leverages the filtration process to provide efficient, low-dimensional summaries rooted in topological intuition.
@@
  • Competitive Performance: Experiments on benchmark graph classification and clustering tasks demonstrate that TopER achieves competitive results compared to more complex, state-of-the-art models, while offering superior interpretability.
  • Interpretable Visualizations: TopER produces low-dimensional outputs that support intuitive visualization of clusters and structural outliers within and across graph datasets, enabling a form of visual model comparison often lacking in traditional embeddings.
  • Robustness Under Perturbations: We provide a stability analysis showing that TopER’s representations are consistent under small changes to the filtration function, reinforcing its suitability for practical and comparative applications.
2 Background
2.1 Related Work
@@
Graph-level Embedding Methods. Graph representation learning, including GNNs and Graph Pooling techniques, is a dynamic subfield of machine learning, focusing on transforming graph data into efficient, low-dimensional vector representations that encapsulate essential features of the data (Ham, 20; GHT+, 19). These representations facilitate a deeper analytical understanding of graphs, which is critical for various applications such as molecular graph property prediction (DTRF, 19). GNNs have revolutionized the analysis of graph data, drawing parallels with the success of Convolutional Neural Networks in image processing (EPBM, 20). GNNs utilize spectral and spatial approaches to graph convolutions based on the graph Laplacian and direct graph convolutions, respectively (BZSL, 14; DBV, 16; KW, 17). Despite their success, GNNs often suffer from issues like over-smoothing and a lack of transparency, making them less ideal for applications requiring interpretability (Gün, 22).
Graph Pooling emerged as a key component in GNN architectures, drawing parallels to the role of pooling in Convolutional Neural Networks  (EPBM, 20). Pooling aims to deduce into meaningful graph embeddings through node aggregation (mean, max, and add pooling  (XHLJ, 19)) or hierarchical pooling (node selection: Top-k  (GJ, 19) and SAGPool  (LLK, 19); node clustering: DiffPool  (YYM+, 18) and MinCutPool  (BGA, 19)). Both pooling groups have their challenges. Node aggregation methods may lose the structural information by treating each node equally, while the hierarchical pooling methods can be computationally heavy, and loss of information can occur if important nodes are discarded.
@@
Topology-inspired Representations. TDA provides a robust and computationally efficient framework to address the interpretability and over-smoothing issues present in GNNs (AAF, 19). Persistent Homology, a key technique in TDA, has been applied successfully to graph data, demonstrating potential to match or even exceed the performance of traditional methods in classification and clustering tasks (HMR, 21; DCG+, 22; HIL+, 24; ISG, 23; CSA+, 23; LSC+, 23). However, the computational intensity of PH limits its scalability (HKN, 19; ZYCW, 20; AKGC, 22).
@@
Visualization Techniques. Graph embedding techniques, including spectral methods, random walk-based approaches, and deep learning-based models, transform graph data into vector representations to support tasks like visualization and machine learning (CZC, 18; GF, 18; Xu, 21). Approaches such as Laplacian Eigenmaps and DeepWalk have been particularly effective in revealing clusters within graphs (BN, 01; PAS, 14). However, these methods are predominantly applied to visualize a single graph in node classification tasks, focusing on cluster identification (WCZ, 16; MK, 20; TPPM, 23). Furthermore, they often overlook domain-specific information, which can limit their effectiveness in more specialized applications (JZ, 20).
@@ 2.2 Persistent Homology for Graphs
Topological Data Analysis (TDA) offers a powerful framework for graph representation learning (AAF, 19), with persistent homology (PH) being especially effective at capturing multi-scale topological features (CA, 24). While PH typically involves filtrations, persistence diagrams, and vectorization, our model focuses on the filtration step, reformulating the evolution of topological features in a novel and efficient way.
In the crucial filtration step, PH decomposes a graph $\mathcal{G}$ into a nested sequence of subgraphs $\mathcal{G}_{1}\subseteq\mathcal{G}_{2}\subseteq\ldots\subseteq\mathcal{G}_{n}=\mathcal{G}$. For each $\mathcal{G}_{i}$, an abstract simplicial complex $\widehat{\mathcal{G}}_{i}$ is defined, forming a filtration of simplicial complexes. Clique complexes are typical choices, where each $(k+1)$-complete subgraph in $\mathcal{G}$ corresponds to a $k$-simplex (AAF, 19).
Filtration. Utilizing relevant filtration functions is essential to obtain effective filtrations. For a given graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, a common approach is to define a node filtration function $f:\mathcal{V}\to\mathbb{R}$, which establishes a hierarchy among the nodes. By selecting a monotone increasing set of thresholds $\mathcal{I}=\{\epsilon_{i}\}_{i=1}^{n}$, this method generates subgraphs $\mathcal{G}_{i}=(\mathcal{V}_{i},\mathcal{E}_{i})$ where $\mathcal{V}_{i}=\{v\in\mathcal{V}\mid f(v)\leq\epsilon_{i}\}$ and $\mathcal{E}_{i}$ is the set of edges in $\mathcal{E}$ with endpoints in $\mathcal{V}_{i}$. This is called a sublevel filtration induced by $f$ (See Figure˜2). Also, superlevel filtrations can be constructed by defining $\mathcal{V}_{i}=\{v\in\mathcal{V}\mid f(v)\geq\epsilon_{i}\}$ for decreasing thresholds (AAF, 19).
Similarly, one can use edge filtration functions $g:\mathcal{E}\to\mathbb{R}$ to define such a filtration. We define $\mathcal{E}_{i}=\{e_{jk}\in\mathcal{E}\mid g(e_{jk})\leq\epsilon_{i}\}$, and $\mathcal{V}_{i}$ as all the endpoints of $\mathcal{E}_{i}$ to create a nested sequence $\{\mathcal{G}_{i}\}_{i=1}^{n}$. Especially for weighted graphs, this method is highly preferable as weights naturally define an edge filtration function. The common node filtration functions are degree, betweenness, centrality, heat kernel signatures (BK, 10), and node functions coming from the domain of the datasets (e.g., atomic number for molecular graphs). Common edge filtration functions are Ollivier and Forman Ricci curvatures (LLY, 11) and edge weights (e.g., transaction amounts for financial networks).
@@
In addition to existing approaches, we introduce a new filtration function, Popularity, which extends the idea of degree-based ranking (New, 03). While the degree function measures the number of direct neighbors a node has, Popularity captures the average degree of those neighbors. The underlying intuition is that, whereas degree reflects how many connections a node has, Popularity emphasizes node influence through association with high-degree nodes.
Formally, for each node $v$, we define popularity as: $\mathcal{P}(v)=\deg(v)+\frac{\sum_{u\in\mathcal{N}(v)}\deg(u)}{|\mathcal{N}(v)|}$ where $\deg(v)$ is the node’s degree and $\mathcal{N}(v)$ is its set of neighbors. This function incorporates 2-neighborhood information by weighting high-degree neighbors more heavily, making it a refined version of the degree function.
@@ 3 TopER: Topological Evolution Rate
TopER is inspired by the foundational idea that topology can capture the shape of graph data, and that this shape can be observed through the evolution of a graph during the filtration process. In this sense, TopER is topology-inspired. However, TopER diverges from traditional PH in how it tracks the shape. We first achieve a computationally efficient alternative to persistent homology by simplifying its filtration-based perspective, and second, we develop a low-dimensional, interpretable representation of graphs that enables both effective classification and intuitive visualization across graph datasets.
Our reformulation reduces the computational overhead typically required for topological feature extraction. Unlike Persistent Homology, which extracts costly topological features, TopER summarizes the filtration process through two key parameters derived via regression: filtration sequences and evolution.
Filtration sequences. We first decompose a graph $\mathcal{G}$ into a nested sequence of subgraphs (filtration graphs) $\mathcal{G}_{1}\subseteq\mathcal{G}_{2}\ldots\subseteq\mathcal{G}_{n}=\mathcal{G}$ by using a filtration function, such as node degree or closeness. Let $\mathcal{G}_{i}\subset\mathcal{G}$, $\mathcal{V}_{i}$ represent nodes in $\mathcal{G}_{i}$ and $\mathcal{E}_{i}$ represent the edges. Next, we compute $x_{i}=|\mathcal{V}_{i}|$ as the count of nodes, and $y_{i}=|\mathcal{E}_{i}|$ as the count of edges. Then, for each filtration graph $\mathcal{G}_{i}$, we obtain the pair $(x_{i},y_{i})\in\mathbb{R}^{2}$, which creates two monotone sequences, $x_{1}\leq x_{2}\leq\dots\leq x_{n}$ and $y_{1}\leq y_{2}\leq\dots\leq y_{n}$. Hence, TopER yields two ordered sets $\mathcal{X},\mathcal{Y}$ describing the evolution of the filtration graphs $\mathcal{G}_{1}\subseteq\ldots\subseteq\mathcal{G}_{n}=\mathcal{G}$, $\mathcal{X}=(x_{1},x_{2},\dots,x_{n})$ and $\mathcal{Y}=(y_{1},y_{2},\dots,y_{n})$. Here, $n$ corresponds to number of thresholds $\{\epsilon_{i}\}_{i=1}^{n}$ used in the filtration step. Consider the top row of Figure˜2 where we have three filtration graphs (i.e., nn=3); we have $\mathcal{X}=(2,3,7)$ for node counts and $\mathcal{Y}=(0,2,6)$ for edge counts.
@@
Evolution. In the next step, PH would typically compute topological features on each filtration and create a persistence diagram to summarize the features. Not only is it costly, but the approach would require efforts to vectorize the persistence diagrams. We circumvent this computationally costly step and analyze how the number of edges $\{y_{i}\}$ relates to the number of nodes $\{x_{i}\}$ throughout the filtration sequence. We use line fitting to characterize this relationship as follows. Simple linear regression, often applied through the least squares method (JWH+, 23), is a standard approach in regression analysis for fitting a linear equation to a set of data points $\{(x_{i},y_{i})\}\subset\mathbb{R}^{2}$. This method calculates the line $L(x)=a+bx$ that best fits the data by minimizing the loss function $\mathrm{E}=\sum^{N}_{i=1}[L(x_{i})-y_{i}]^{2}$. The regression coefficients $(a,b)$ capture the graph’s structural evolution through filtration (see descriptor step in Figure˜3). The full TopER method is outlined in Algorithm˜1 in the Appendix.
@@
Definition 3.1 (Topological Evolution Rate (TopER)).
Let $f:\mathcal{V}\to\mathbb{R}$ be a filtration function on graph $\mathcal{G}$ and $\mathcal{I}=\{\epsilon_{i}\}_{i=1}^{n}$ be the threshold set. Let $\mathcal{G}_{i}=(\mathcal{V}_{i},\mathcal{E}_{i})$ be the induced filtration. Let $x_{i}=|\mathcal{V}_{i}|$ and $y_{i}=|\mathcal{E}_{i}|$. Let $L(x)=a+bx$ be the best fitting line to $\{(x_{i},y_{i})\}_{i=1}^{n}$. Then, we define the TopER vector of $\mathcal{G}$ wrt. $f$ as $\rm{TE}_{f}(\mathcal{G},\mathcal{I})=(a,b)$. We call $a$ the pivot and $b$ the growth of $\mathcal{G}$.
Remark 3.2 (Why a Linear Fit?).
Although one could consider higher-order polynomials, our experiments show that coefficients beyond the first degree are negligible (Table˜15), revealing an essentially linear relationship. Due to space constraints, we defer a more detailed comparison of linear, polynomial, and higher-order fits to Appendix˜B, where we also provide visual examples of evolution rates (Figure˜6) and fundamental pattern types (Figure˜7).
Remark 3.3 (On the Name TopER).
Although TopER does not compute classical topological invariants such as persistence diagrams or Betti numbers, its name is grounded in two topology-inspired principles. First, it mirrors the TDA notion of a filtration, a nested sequence of subgraphs $\{\mathcal{G}_{i}\}$ that progressively unveils structural features across scales. Second, it is directly tied to the evolution of the Euler characteristic $\chi(\mathcal{G}_{i})\;=\;|\mathcal{V}_{i}|-|\mathcal{E}_{i}|\;=\;\beta_{0}(\mathcal{G}_{i})-\beta_{1}(\mathcal{G}_{i}),$ where $\beta_{i}$ denotes the $i$th Betti number. Equivalently, one could track $\{(x_{i},y_{i}-x_{i})\}=\{(|\mathcal{V}_{i}|,\chi(\mathcal{G}_{i}))\}$, but we use $\{(x_{i},y_{i})\}$ for notational simplicity. The resulting slope thus quantifies how topological complexity changes over the filtration, justifying the term Topological Evolution Rate.
We emphasize that the term “topological” reflects the conceptual roots of TopER in topological data analysis: the method uses a filtration to reveal structural patterns in the graph, analogous to the filtration process in persistent homology. While we do not compute full persistence diagrams or Betti numbers, the slope of the Euler characteristic across the filtration captures essential topological information, justifying the terminology.